# Probing the Effective Quantum Gravity via Quasinormal Modes and Shadows of Black Holes

### v1

@@ Abstract
Two quantum-corrected black hole models have recently been proposed within the Hamiltonian constraints approach to quantum gravity, maintaining general covariance Zhang:2024khj . We have studied the quasinormal spectra of these black holes using three methods: the higher-order WKB approach with Padé approximants, time-domain integration, and the Frobenius method. The Frobenius method, in particular, allows us to determine precise values of the frequencies, including the overtones. The two models differ in their choice of quantum parameter $\xi$, and we can distinguish them by their quasinormal spectra. In the first model, increasing the quantum parameter results in higher real oscillation frequencies and damping rates of the fundamental mode. In contrast, the second model shows a decrease in the oscillation frequency of the least-damped mode when the quantum parameter is introduced. We have shown that, while the fundamental mode changes relatively gradually with the quantum parameter, the first few overtones deviate from their Schwarzschild limits at an increasing rate. This results in a qualitatively new behavior: the real parts of the frequencies of the first and higher overtones tend to zero as the quantum parameter increases. In addition to the branch of modes that are perturbative in the quantum parameter, we observe some non-perturbative modes at moderate values of the quantum parameter. Additionally, we have calculated the radii of the shadows cast by these black holes and discussed possible constraints based on observations of Sgt $A^{*}$.
@@ I Introduction
Recently, a long-standing issue regarding general covariance in spherically symmetric gravity, which arises when canonical quantum gravity leads to a semiclassical model of black holes, was addressed in Zhang:2024khj . This work proposed two black hole models that differ based on the choice of a quantum parameter. While numerous studies have explored quasinormal modes of quantum-corrected black holes Bolokhov:2023bwm ; Bolokhov:2023ruj ; Skvortsova:2024atk ; Dubinsky:2024fvi ; Dubinsky:2024aeu ; Bolokhov:2023ozp ; Konoplya:2019xmn ; Konoplya:2023ahd ; Gong:2023ghh ; Cruz:2020emz ; Daghigh:2020fmw ; 2024arXiv240707892S ; 2024arXiv240719654Z , the models produced within the effective Hamiltonian approach, maintaining general covariance, as presented in Zhang:2024khj , have not been previously considered.
@@
In this study, we investigate the quasinormal frequencies of the two black hole metrics derived in Zhang:2024khj for scalar, electromagnetic, and effective axial gravitational perturbations. We employ three independent methods for these calculations: the higher-order WKB approach with Padé approximants, time-domain integration, and the Frobenius method. The Frobenius method, in particular, provides precise results, enabling us to determine accurate values for the overtones. Our findings reveal that while the fundamental mode changes relatively mildly upon the introduction of quantum corrections, the first few overtones deviate significantly, with their real oscillation frequencies tending towards zero. This behavior creates a distinctive "sound" of the event horizon deformed by quantum corrections. When the quantum parameter is not sufficiently small, we observe quasinormal modes that do not transition into the Schwarzschild modes as the quantum parameter approaches zero.
@@
The structure of our work is as follows. In Section II, we introduce the black hole metrics and discuss the wave equations for scalar, electromagnetic, and effective gravitational perturbations. Section III reviews the three methods used for calculating the quasinormal frequencies. In Section IV, we present the numerical results for the quasinormal modes and provide analytical solutions in the eikonal regime. Finally, in the Conclusions, we discuss our findings and highlight some open problems.
@@ II.1 The metric and the underlying theory
The advantage of this black hole metric is that it maintains the same relationship between the event horizon radius $r_{0}$ and the black hole mass as the classical solution, $r_{0}=2M$. This consistency facilitates an easy comparison between the spectra of classical and quantum-corrected black holes.
@@ II.2 Perturbation of test fields
The effective potentials for the above fields are shown in figs. 1-3. They are positive definite, which guarantees stability for these perturbations.
@@ II.3 Perturbations of gravitational field
$h_{1}=\frac{r}{\sqrt{f(r)g(r)}}\Psi,\quad dr^{*}=\frac{dr}{\sqrt{f(r)g(r)}},$ (18)
@@
$\frac{d^{2}}{dr_{*}{}^{2}}\Psi+(\omega^{2}-V_{ax}(r)=0,$ (19)
@@
Figure 2: Effective potential as a function of the tortoise coordinate of the $\ell=1$ sscalar field perturbations of the first black hole model ($M=1/2$): $\xi=0$ (black), $\xi=0.4$ (blue), $\xi=0.8$ (green), $\xi=1.2$ (orange).
III WKB method and time-domain integration
Here we briefly review the three methods used for the calculations of quasinormal modes: the WKB method, time-domain integration, and the Frobenius method. The latter method is based on a converging procedure and, therefore, provides precise values of quasinormal frequencies. By definition, quasinormal modes satisfy the following boundary conditions,
@@ III.3 Frobenius method
such that the factor $P(r,\omega)$ makes $y(r)$ regular in the range $r_{0}\leq r$ when the quasinormal modes boundary conditions are satisfied. Then, expanding $y(r)$ as follows
@@
$P(r,\omega)=\left(\frac{r-r_{h}^{+}}{r-r_{h}^{-}}\right)^{-\frac{i\omega}{f^{\prime}(r_{h}^{+})}}(r-r_{h}^{-})^{2i\omega r_{h}^{+}}e^{i\omega r},$ (28)
@@
Figure 6: The fundamental mode and first four overtones as a function of $\xi$ for the first black-hole model; $\ell=0$ scalar field perturbations.
@@
Figure 7: The fundamental mode and first four overtones as a function of $\xi$ for the first black-hole model; $\ell=1$ electromagnetic field perturbations.
@@
Figure 8: The fundamental mode and first four overtones as a function of $\xi$ for the second black-hole model; $\ell=0$ scalar field perturbations.
@@
Figure 9: The fundamental mode and first four overtones as a function of $\xi$ for the second black-hole model; $\ell=2$ axial gravitational perturbations.
@@
Figure 10: The fundamental mode and first five overtones as a function of $\xi$ for the second black-hole model; $\ell=2$ axial gravitational perturbations.
@@
Figure 11: The perturbative branch (black) of the third overtone and two non-pertubative modes (blue, red) as a function of $\xi$ for the first black-hole model; $\ell=2$ axial gravitational perturbations.
@@ IV Quasinormal modes
Figure 12: Time-domain profiel for axial gravitational perturbations of the first black hole model, $\ell=2$, $M=1/2$, $\xi=0.3$.
@@
The fundamental quasinormal modes for both black hole models are presented in Tables 1-10. A notable difference between the modes of the first and second models is observed in the real oscillation frequency. For the second model, this frequency changes very slightly as the parameter $\xi$ is introduced, whereas it increases with $\xi$ in the first model. This behavior can be attributed to the differences in the black hole metrics: in the first model, both the components $g_{tt}$ and $g_{rr}$ differ from the Schwarzschild case, while in the second model, $g_{tt}$ retains the Schwarzschild form. As a result, the real part of the frequency, $\text{Re}(\omega)$, which is primarily determined by the centrifugal part of the effective potential, exhibits only minor changes with $\xi$.
It can be observed that the real oscillation frequency of the fundamental mode increases monotonically with the parameter $\xi$ for perturbations of all types of the first model and for scalar perturbations of the second model. However, this is not the case of the electromagnetic and gravitational perturbations of the second black hole model. Such behavior could be explained by the form of the effective potentials which are higher for larger values of $\xi$ (see figs. 1-4 ), while this is not so for the perturbations of the second model, as can be seen directly from the analytic expression for the potential 5. The damping rate also generally increases with $\xi$, up to a certain relatively large value, beyond which it may start to decrease, as illustrated in Fig. 7. A comparison of the WKB approach, time-domain integration, and the precise results of the Frobenius method indicates that while the WKB method can be more accurate for $\ell=0$, the time-domain integration becomes more reliable for larger values of $\ell$. This difference in accuracy arises because, at $\ell=0$, the ringing period is very short, making it challenging for the Prony method to extract frequencies with high precision.
@@
This relation, which connects the grey-body factors $\Gamma_{\ell}(\omega)$ with the fundamental quasinormal mode $\omega_{0}$, is exact in the eikonal limit as $\ell\to\infty$. By introducing corrections involving the first overtone $\omega_{1}$, this relation also provides an approximate correspondence for lower $\ell$. Consequently, the quasinormal modes obtained here can be used to determine the grey-body factors. In the eikonal regime, the grey-body factors can be immediately derived analytically using the eikonal expressions (30) and (31) for the quasinormal modes $\omega_{1}$ and $\omega_{2}$ in (32).

### v3

@@ Abstract
Two quantum-corrected black hole models have recently been proposed within the Hamiltonian constraints approach to quantum gravity, maintaining general covariance Zhang:2024khj . We have studied the quasinormal spectra of these black holes using four methods: the higher-order WKB approach with Padé approximants, time-domain integration, Frobenius, and pseudospectral methods. The Frobenius method, in particular, allows us to determine precise values of the frequencies, including the overtones. The two models differ in their choice of quantum parameter $\xi$, and we can distinguish them by their quasinormal spectra. In the first model, increasing the quantum parameter results in higher real oscillation frequencies and damping rates of the fundamental mode. In contrast, the second model shows a decrease in the oscillation frequency of the least-damped mode when the quantum parameter is introduced. We have shown that, while the fundamental mode changes relatively gradually with the quantum parameter, the first few overtones deviate from their Schwarzschild limits at an increasing rate. This results in a qualitatively new behavior: the real parts of the frequencies of the first and higher overtones tend to zero as the quantum parameter increases. In addition to the branch of modes that are perturbative in the quantum parameter, we observe some non-perturbative modes at moderate values of the quantum parameter. Additionally, we have calculated the radii of the shadows cast by these black holes and discussed possible constraints based on observations of Sgt $A^{*}$. As a byproduct, we tested the method of calculating quasinormal modes of this kind based on a recent parametrization of effective potentials, and showed that while the parametrized formalism could be used for estimating the fundamental mode at small values of the coupling, its accuracy is highly dependent on the particular spacetime under consideration and is insufficient even for the lowest overtones.
@@ I Introduction
Recently, a long-standing issue regarding general covariance in spherically symmetric gravity, which arises when canonical quantum gravity leads to a semiclassical model of black holes, was addressed in Zhang:2024khj . This work proposed two black hole models that differ based on the choice of a quantum parameter.
The fundamental characteristic of black hole geometry is its spectrum of quasinormal modes Kokkotas:1999bd ; Konoplya:2011qq , which can be observed through gravitational wave interferometers LIGOScientific:2017vwq ; LIGOScientific:2020zkf . Future experiments promise to detect a much broader range of frequencies LISA:2022kgy . While the fundamental mode primarily depends on the peak of the potential barrier, the first few overtones describe the geometry near the event horizon Konoplya:2023hqb ; Konoplya:2022pbc . In the time domain, these first overtones are crucial for describing the initial stage of the ringdown Giesler:2019uxc . Observations in the gravitational spectrum can be complemented by electromagnetic observations, such as measuring the shadows cast by black holes EventHorizonTelescope:2019dse and analyzing other optical phenomena Bambi:2015kza .
While numerous studies have explored quasinormal modes of quantum-corrected black holes Bolokhov:2023bwm ; Bolokhov:2023ruj ; Skvortsova:2024atk ; Dubinsky:2024fvi ; Dubinsky:2024aeu ; Bolokhov:2023ozp ; Konoplya:2019xmn ; Konoplya:2023ahd ; Gong:2023ghh ; Cruz:2020emz ; Daghigh:2020fmw ; 2024arXiv240707892S ; 2024arXiv240719654Z ; Calza:2024fzo ; Calza:2024xdh , the recent models produced within the effective Hamiltonian approach, maintaining general covariance, as presented in Zhang:2024khj , have not been previously considered.
@@
In this study, we investigate the quasinormal frequencies of the two black hole metrics derived in Zhang:2024khj for scalar, electromagnetic, and effective axial gravitational perturbations. We employ four independent methods for these calculations: the higher-order WKB approach with Padé approximants, time-domain integration, the Frobenius method and the pseudospectral method. The Frobenius method, in particular, provides precise results, enabling us to determine accurate values for the overtones. Our findings reveal that while the fundamental mode changes relatively mildly upon the introduction of quantum corrections, the first few overtones deviate significantly, with their real oscillation frequencies tending towards zero. This behavior creates a distinctive "sound" of the event horizon deformed by quantum corrections. When the quantum parameter is not sufficiently small, we observe quasinormal modes that do not transition into the Schwarzschild modes as the quantum parameter approaches zero. In addition, we calculate the radius of the shadow cast by these two black holes.
We also used this study as an opportunity to test the parametrization of the effective potential suggested in Cardoso_2019 as a tool for finding quasinormal modes. For this purpose, we have considered not only the two quantum-corrected black hole models developed in Zhang:2024khj , but also the black hole spacetime recently found as a result of quantum-corrected collapse in  Lewandowski_2023 . While test field perturbations have been recently considered for this model in Gong:2023ghh ; Skvortsova:2024atk ; Zinhailo:2024kbq ; Luo:2024dxl , no analysis of gravitational spectra has been done so far. In this way, we also complemented the existing works by studying the most valuable gravitational perturbations. It turns out that the parametrized approach Cardoso_2019 can be used for calculating the fundamental quasinormal modes only, and even in that case, its accuracy is highly dependent on the model under consideration. For overtones, the relative error is usually of the same order as (or larger than) the effect (i.e., the deviation of the frequencies from their Schwarzschild values).
@@
The structure of our work is as follows. In Section II, we introduce the black hole metrics and discuss the wave equations for scalar, electromagnetic, and effective gravitational perturbations. Section III reviews the four methods used for calculating the quasinormal frequencies. In Section IV, we present the numerical results for the quasinormal modes and provide analytical solutions in the eikonal regime. Section V is devoted to the discussion of the shadows cast by these black holes. Finally, in the Conclusions, we discuss our findings and highlight some open problems.
@@ II.1 The metric and the underlying theory
The advantage of this black hole metric is that it maintains the same relationship between the event horizon radius $r_{h}^{+}$ and the black hole mass as the classical solution, $r_{h}^{+}=2M$. This consistency facilitates an easy comparison between the spectra of classical and quantum-corrected black holes.
@@ II.2 Perturbation of test fields
The effective potentials for the above fields are shown in figs. 13. They are positive definite, which guarantees stability for these perturbations.
@@ II.3 Perturbations of gravitational field
Figure 2: Effective potential as a function of the tortoise coordinate of the $\ell=1$ scalar field perturbations of the first black hole model ($M=1/2$): $\xi=0$ (black), $\xi=0.4$ (blue), $\xi=0.8$ (green), $\xi=1.2$ (orange).
@@
$h_{1}=\frac{r}{\sqrt{f(r)g(r)}}\Psi,\quad dr_{*}=\frac{dr}{\sqrt{f(r)g(r)}},$ (18)
@@
$\frac{d^{2}}{dr_{*}^{2}}\Psi+\left[\omega^{2}-V_{ax}(r)\right]=0,$ (19)
@@
III Methods
Here we briefly review the four methods used for the calculations of quasinormal modes: the WKB method, time-domain integration, the Frobenius method, and the pseudospectral method. The Frobenius method is based on a converging procedure and, therefore, provides precise values of quasinormal frequencies.
By definition, quasinormal modes satisfy the following boundary conditions,
@@ III.3 Frobenius method
such that the factor $P(r,\omega)$ makes $y(r)$ regular in the range $r_{h}^{+}\leq r$ when the quasinormal modes boundary conditions are satisfied. Then, expanding $y(r)$ as follows
@@
$P(r,\omega)=\left(\frac{r-r_{h}^{+}}{r-r_{h}^{-}}\right)^{-\frac{i\omega}{f^{\prime}(r_{h}^{+})}}(r-r_{h}^{-})^{2ir_{h}^{+}\omega}e^{i\omega r},$ (28)
@@
III.4 Pseudospectral method
The Chebyshev pseudospectral method is a powerful technique for solving differential equations. The method is based on discretizing the unknown function over a grid of collocation points, typically corresponding to the roots of Chebyshev polynomials. By transforming the differential equations into a system of algebraic equations at these collocation points and solving them, we can determine the function’s values at grid points.
To ensure the correct boundary conditions corresponding to quasinormal modes, we make a substitution in the master equation (3)
$\Psi(r)=r^{2ir_{h}^{+}\omega}(r-r_{h}^{+})^{-ir_{h}^{+}\omega}e^{i\omega r}y(r),$ (29)
which leads to $y\sim const$ at the outer black hole horizon and at spatial infinity.
Then, we compactify our semi-interval $[r_{h},\infty)$ to interval $[0,1]$ by introducing a new variable
$r=\frac{r_{h}}{1-u},$ (30)
Altogether, these substitutions lead to an equation of the form
$A_{2}(u)y^{\prime\prime}(u)+A_{1}(u)y^{\prime}(u)+A_{0}(u)y(u)=0,$ (31)
where $A_{i}(u)=A_{i}(u,\omega,\omega^{2})$, $i=0,1,2$. At the next step, we discretize equation (31) on the Chebyshev-Lobatto grid, which is defined as
$u_{j}=\frac{1}{2}\left(1-\cos\left[\frac{\pi j}{N}\right]\right),~{}~{}j=0,1...N.$ (32)
This results in a matrix equation
$\left(\tilde{M}_{0}+\tilde{M}_{1}\omega+\tilde{M}_{2}\,\omega^{2}\right)\tilde{y}=0,$ (33)
where $\tilde{y}$ is the vector of the unknown function’s values, $\tilde{M}_{i}$ are the numerical matrices of discretized coefficients at the collocation grid points.
We can linearize our quadratic eigenproblem (33) in the following way
$\left(M_{0}+M_{1}\omega\right)\tilde{\psi}=\mathbb{0},$ (34)
where
$M_{0}=\begin{pmatrix}\tilde{M}_{0}&\tilde{M}_{1}\\ \mathbb{0}&\mathbb{1}\end{pmatrix},\,M_{1}=\begin{pmatrix}\mathbb{0}&\tilde{M}_{2}\\ -\mathbb{1}&\mathbb{0}\end{pmatrix},\,\tilde{\psi}=\begin{pmatrix}\tilde{y}\\ \omega\tilde{y}\end{pmatrix}.$ (35)
Then QNMs spectrum can be found by solving generalized eigenvalue problem (34) via the Eigenvalues command in Mathematica. To avoid spurious eigenvalues we perform the calculations on two grids of different sizes and select only overlapping values boyd2013chebyshev .
@@ III.4 Pseudospectral method
Figure 6: The fundamental mode and the first four overtones as a function of $\xi$ for the first black-hole model; $\ell=0$ scalar field perturbations.
@@
Figure 7: The fundamental mode and the first four overtones as a function of $\xi$ for the first black-hole model; $\ell=1$ electromagnetic field perturbations.
@@
Figure 8: The fundamental mode and the first four overtones as a function of $\xi$ for the second black-hole model; $\ell=0$ scalar field perturbations.
@@
Figure 9: The fundamental mode and the first four overtones as a function of $\xi$ for the first black-hole model; $\ell=2$ axial gravitational perturbations.
@@
Figure 10: The fundamental mode and the first five overtones as a function of $\xi$ for the second black hole model; $\ell=2$ axial gravitational perturbations.
@@
Figure 11: The first non-perturbative modes as a function of $\xi$ for the first black hole model; $\ell=2$ axial gravitational perturbations. The star symbols correspond to the values from the Table. 9
Figure 12: The complex $\omega$ plane of QNMs. : The black lines represent the perturbative branches, while the colored lines correspond to the non-perturbative branches for the first black hole model; $\ell=2$ axial gravitational perturbations.
Figure 13: Time-domain profile for axial gravitational perturbations of the first black hole model, $\ell=2$, $M=1/2$, $\xi=0.3$.
@@
$\xi$ PS Frobenius method rel. diff. $Re(\omega)$ rel. diff. $Im(\omega)$
$1$ $0.102-3.0901i$ $0.102225-3.090315i$ $0.157\%$ $0.012\%$
$1.13$ $0.1212-2.5679i$ $0.0904-3.1989i$ $0.121201-2.567954i$ $0.08986-3.19839i$ $0.0095\%$ $0.562\%$ $0.0025\%$ $0.017\%$
$1.15$ $0.05503-1.9585i$ $0.124-2.5799i$ $0.053611-1.9592i$ $0.124053-2.579967i$ $2.65\%$ $0.0378\%$ $0.0384\%$ $0.001\%$
$1.2$ $0.11921-1.973i$ $0.126-2.6106i$ $0.119201-1.973059i$ $0.125993-2.61059i$ $0.0048\%$ $0.011\%$ $0.001\%$ $0.0019\%$
$2$ $0.0989-2.6595i$ $0.099091-2.659618i$ $0.1711\%$ $0.0037\%$
Table 9: Comparison of the nonperturbative quasinormal frequencies for axial gravitational perturbations ($s=2$) of the first BH model obtained by the pseudospectral (PS) method and by the Frobenius method: $\ell=2$, $M=1/2$.
@@ IV Quasinormal modes
The fundamental quasinormal modes for both black hole models are presented in Tables 1-8. A notable difference between the modes of the first and second models is observed in the real oscillation frequency. For the second model, this frequency changes very slightly as the parameter $\xi$ is introduced, whereas it increases with $\xi$ in the first model. This behavior can be attributed to the differences in the black hole metrics: in the first model, both the components $g_{tt}$ and $g_{rr}$ differ from the Schwarzschild case, while in the second model, $g_{tt}$ retains the Schwarzschild form. As a result, the real part of the frequency, $\text{Re}(\omega)$, which is primarily determined by the centrifugal part of the effective potential, exhibits only minor changes with $\xi$.
It can be observed that the real oscillation frequency of the fundamental mode increases monotonically with the parameter $\xi$ for perturbations of all types of the first model and for scalar perturbations of the second model. However, this is not the case of the electromagnetic and gravitational perturbations of the second black hole model. Such behavior could be explained by the form of the effective potentials which are higher for larger values of $\xi$ (see figs. 1-4 ), while this is not so for the perturbations of the second model, as can be seen directly from the analytic expression for the potential fig. 5. The damping rate also generally increases with $\xi$, up to a certain relatively large value, beyond which it may start to decrease, as illustrated in Fig. 7. A comparison of the WKB approach, time-domain integration, and the precise results of the Frobenius method indicates that while the WKB method can be more accurate for $\ell=0$, the time-domain integration becomes more reliable for larger values of $\ell$. This difference in accuracy arises because, at $\ell=0$, the ringing period is very short, making it challenging for the Prony method to extract frequencies with high precision.
@@
This relation, which connects the grey-body factors $\Gamma_{\ell}(\omega)$ with the fundamental quasinormal mode $\omega_{0}$, is exact in the eikonal limit as $\ell\to\infty$. By introducing corrections involving the first overtone $\omega_{1}$, this relation also provides an approximate correspondence for lower $\ell$. Consequently, the quasinormal modes obtained here can be used to determine the grey-body factors. In the eikonal regime, the grey-body factors can be immediately derived analytically using the eikonal expressions (37) and (38) for the quasinormal modes $\omega_{1}$ and $\omega_{2}$ in (39). This correspondence also applies to a variety of rotating black holes Konoplya:2024vuj .